### v2

# Sharp transitions in small exciton spectra for multi-orbital lattice systems

Abstract
We demonstrate that strongly bound excitons, whose radii approach the lattice constant, display physics that eludes continuum descriptions not just quantitatively but qualitatively. We investigate such phenomenology by calculating the exciton spectra for several one- and two-dimensional lattice models, whose valence band has a multi-orbital character. We identify sharp transitions in the character and momentum of the lowest-energy exciton, driven directly by the multi-orbital nature of the lattice models. Such transitions cannot occur in simple continuum descriptions.
preprint: APS/123-QED
I Introduction
Excitons—electron-hole pairs bound by their mutual Coulomb attraction—are observed in the optical response of most semiconductors. In conventional inorganic semiconductors such as Si, the strength of the attraction is small compared to the bandwidths of the conduction and valence bands, and the resulting Wannier-Mott excitons [1, 2, 3, 4] are weakly bound and have a radius that greatly exceeds the lattice constant $a$. Consequently, a continuum description suffices: the maximum and minimum of valence and conduction bands are well approximated as quadratic, describing well the exciton wavefunction which is dominated by the contribution of the optima of these bands. For three-dimensional (3D) crystals and a $1/r$ Coulomb attraction, this maps the exciton spectrum to that of a hydrogen atom, with a rescaled Bohr radius and Rydberg energy [2, 3, 4, 5, 6].
For a screened, short-range attraction with a characteristic scale $U$, excitons appear in 3D only if $U$ exceeds a critical value. However, both in two dimensions (2D) and one dimension (1D), a bound state exists for arbitrarily small $U$, with a binding energy $E_{\text{bind}}\propto\exp(-W/U)$ and $E_{\text{bind}}\propto U^{2}/W$, respectively [5, 8, 7] (here $W$ is a measure of the average kinetic energy of the pair). These asymptotic estimates are widely used in standard many-body approaches to excitons [9, 10] and provide a baseline for comparison between theory and experiments, for example, see Refs. 11, 12.
However, many low-dimensional semiconductors that are technologically relevant, host strongly bound Frenkel and charge-transfer excitons [13, 14, 15] whose radii are equal or comparable to the lattice constant $a$. For example, TR-ARPES on epitaxial C60 films on Au(111) has revealed binding energies of 300–1000 meV for excitons that are either on the same C60 molecule, or have the electron and hole on neighbor molecules [17, 16, 18, 19, 20]. These results agree with earlier work on C60 using other optical techniques [21, 22]. Other examples include organic crystals such as pentacene and PTCDA, which also exhibit exciton binding energies on the order of 0.5 eV [23]. In monolayer WS2, two-photon spectroscopy has measured an exciton binding energy of $\sim 0.3$ eV, underlining that atomically thin semiconductors host non-hydrogenic excitons [24]. For these small excitons, the specific lattice symmetry becomes relevant beyond just the renormalization of the conduction and valence band effective masses [25]. This is because the strong attraction mixes within the exciton wavefunction hole/electron states lying far below/above the top/bottom of the valence/conduction band, and their dispersion is not described well by the continuum approximation. Furthermore, especially in organic semiconductors, the conduction and valence bands often have multi-orbital character, leading to distinct spectral fingerprints and momentum-dependent symmetry changes [26, 27, 28] which may be completely missed by a continuum approximation.
These considerations show that the convenient, simple estimates provided by the continuum approximation must become quantitatively inaccurate once small excitons are concerned, yet it is generally believed that they remain qualitatively correct. In this work, we show that it is very easy to find examples where these simple estimates are qualitatively wrong, especially for small excitons in multi-orbital models. Fortunately, solving such two-particle problems (here, the electron and hole) on an infinite lattice can be done very efficiently, as we demonstrate with a method adapted from studies of few-particle interacting systems, see e.g. Ref. 29. We hope that such methods will be adopted to study realistic lattice models and obtain accurate properties of their small excitons.
The article is organized as follows: Section II introduces the lattice models we study, and Section III details our formalism. Our results are shown in Section IV, and we summarize our findings in Section V.
II The lattice models
The generic Hamiltonian describing an exciton is:
${\cal H}={\cal H}_{e}+{\cal H}_{h}+V_{\rm e-h}$ (1)
where the terms describe the electron moving in the otherwise empty conduction band, the hole moving in the otherwise full valence band, and their Coulomb attraction, respectively.
In our lattice models, we allow the valence band to be multi-orbital:
${\cal H}_{h}=-\sum_{\langle n,m\rangle,\alpha\beta}t_{\alpha\beta}({\bf R}_{n}-{\bf R_{m}})(v^{\dagger}_{\alpha n}v_{\beta m}+H.c.)$ (2)
Here $n,m$ index lattice sites located at ${\bf R}_{n},{\bf R_{m}}$. $\alpha,\beta$ label the specific valence orbitals, and $v^{\dagger}_{\alpha n}$ creates an electron in the orbital $\alpha$ at site $n$. Explicitly, we will show results for 1D models with 2 orbitals per site ($1s$ and $2p_{x}$ for a chain parallel to the $x$ axis) and 2D triangular models with 3 orbitals per site ($d_{x^{2}-y^{2}},d_{xy},d_{3z^{2}-r^{2}}$ for a lattice lying in the $xy$ plane). Any other combinations of orbitals can be studied similarly. Crystal fields can be included trivially, but we ignore them for simplicity. We restrict the hopping to nearest-neighbor for simplicity, but this can be generalized straightforwardly to longer range. To minimize the number of free parameters for the 2D system, we use the Harrison rules [30] to calculate ratios between the hopping integrals, see Appendix A for details. For the same reasons, we present results only for a single-orbital conduction band:
${\cal H}_{e}=-t_{c}\sum_{\langle n,m\rangle}(c^{\dagger}_{n}c_{m}+H.c.)+\Delta\sum_{n}c^{\dagger}_{n}c_{n}$ (3)
where $c_{n}^{\dagger}$ creates a conduction-band electron at site $n$ of the lattice. Here $\Delta$ is the difference between the on-site energies, and controls the value of the band gap. For simplicity we consider again only nearest-neighbor hopping; generalization to longer range hopping is straightforward. Similarly, generalizations to a multi-orbital conduction band is straightforward but introduces more free parameters. We note that in this work we ignore possible hybridization between the orbitals defining the conduction and valence bands. These can be included but lead to more significant complications, in particular lifting of the degeneracy between singlet and triplet excitons. In their absence, the spin degrees of freedom are trivial, thus we drop the spin labels in this work.
Hamiltonians (2) and (3) can be diagonalized straightforwardly to:
${\cal H}_{h}=\sum_{{\bf k},\gamma}E_{v}({\bf k},\gamma)v^{\dagger}_{{\bf k},\gamma}v_{{\bf k},\gamma}$ (4)
where $\bf k$} is the momentum restricted to the Brillouin zone and $\gamma$ indexes the valence bands of energy $E_{v}({\bf k},\gamma)$, and
${\cal H}_{e}=\sum_{{\bf k}}E_{c}({\bf k})c^{\dagger}_{{\bf k}}c_{{\bf k}}$ (5)
defines the conduction band energy $E_{c}({\bf k})$.
We describe the electron-hole attraction as:
$V_{\rm e-h}=-\sum_{n,\alpha}U_{\alpha}c^{\dagger}_{n}c_{n}v_{\alpha,n}v^{\dagger}_{\alpha,n}$ (6)
At first sight, this on-site attraction is an unreasonable approximation for organic semiconductors whose small excitons are due precisely to the very weakly screened, long-range Coulomb interaction. Nevertheless, in the limit of on-site Frenkel excitons of most interest to us in this work, the electron and hole are on the same site with extremely high probability. Because they never 'explore' longer separations, the details of the longer-range Coulomb attraction become irrelevant. The only meaningful interaction energy scale in this limit is the difference between the attraction between an on-site pair and one occupying adjacent sites. This is what the $U_{\alpha}$ energies describe. For completeness, we mention that longer range Coulomb attraction can also be studied with the method we discuss below. A sample calculation for the effect of the nearest neighbor interaction can be found in Appendix C.
III The method
Because of the invariance to translations of the Hamiltonian (1), the exciton momentum $\bf K$} is a good quantum number; its value is restricted to the Brillouin zone of the appropriate lattice. For any value of $\bf K$}, a particularly convenient basis for calculating exciton eigenstates is the set of all states [2, 9, 29]:
$|{\alpha,{\bf K},\bm{\delta}\rangle=\sum_{n}e^{i{\bf K}({\bf R}_{n}+\bm{\delta}/2)}c^{\dagger}_{n}v_{\alpha,n+\bm{\delta}}|{\rm GS}\rangle}$ (7)
where $|{\rm GS}\rangle$ is the ground state with filled valence bands and empty conduction band. Instead of using this basis to solve the Schrödinger equation, it is more convenient to calculate instead the particle-hole propagators [29]:
$G_{\alpha,\beta}({\bf K},\bm{\delta},z)=\langle\alpha,{\bf K},\bm{0}|\hat{G}(z)|\beta,{\bf K},\bm{\delta}\rangle$ (8)
where $\hat{G}(z)=(z-H)^{-1}$ is the resolvent, $z=\omega+i\eta$ is the energy plus a broadening $\eta\rightarrow 0$ and we set $\hbar=1$ and $\eta=10^{-2}t_{c}$ throughout unless otherwise specified. These propagators are useful because according to the Lehmann representation [31, 32]:
$G_{\alpha,\beta}({\bf K},\bm{\delta},z)=\sum_{i}\frac{\langle\alpha,{\bf K},\bm{0}|{\bf K},i\rangle\langle{\bf K},i|\beta,{\bf K},\bm{\delta}\rangle}{z-E({\bf K},i)}$ (9)
[…]
In the absence of Coulomb interactions, these electron-hole eigenstates are $c^{\dagger}_{\bf K+k}v_{{\bf k},\gamma}|{\rm GS}\rangle$ for any $\bf k$}, with the corresponding energy $E_{c}({\bf K+k})-E_{v}({\bf k},\gamma)$. They define the electron-hole continuum that for a given $\bf K$} starts at
$E_{\rm gap}({\bf K})=\min_{{\bf k},\gamma}[E_{c}({\bf K+k})-E_{v}({\bf k},\gamma)]$ (10)
The semiconductor gap is then $E_{gap}=\min_{\bf K}E_{\rm gap}({\bf K})$, and it is a direct gap if the minimum is reached at ${\bf K}=0$; otherwise, it is an indirect gap.
If the Coulomb attraction is sufficiently strong to bind excitons (which is always the case in 1D and 2D systems), their eigenenergies are signalled by discrete poles of $G_{\alpha,\beta}({\bf K},\bm{\delta},z)$ at energies inside the gap $\omega=E_{\rm exc}({\bf K})<E_{\rm gap}({\bf K})$. The corresponding exciton binding energy is $E_{\rm bind}({\bf K})=E_{\rm gap}({\bf K})-E_{\rm exc}({\bf K})$ [4, 32].
Furthermore, the residues associated with each such pole give direct access to that exciton's wavefunction in real space, because
$G_{\alpha,\beta}({\bf K},\bm{\delta},z)\big|_{\omega\rightarrow E_{\rm exc}({\bf K})}\approx\frac{\langle\alpha,{\bf K},\bm{0}|{\bf K},{\rm exc}\rangle\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle}{z-E_{\rm exc}({\bf K})}$ (11)
meaning that
$\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle\propto G_{\alpha,\beta}({\bf K},\bm{\delta},E_{\rm exc}({\bf K})).$ (12)
Eq. (12) gives the real-space exciton amplitude, which we visualize after normalization with the condition $\sum_{\beta,\bm{\delta}}|\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle|^{2}=1$ (Fig. 7,8). We also use it to extract a characteristic exciton radius $\xi$ from the long-range envelope of the eigenstate $\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle\propto\exp(-|\bm{\delta}|/\xi)$ when $|\bm{\delta}|\rightarrow\infty$ [33].
These propagators are calculated from matrix elements between states of Eq. (7) of the identity $\hat{G}(z)(z-{\cal H})=1$, from which we find:
$zG_{\alpha,\beta}({\bf K},\bm{\delta},z)=\delta_{\alpha\beta}\delta_{\bm{\delta},\bm{0}}+\langle\alpha,{\bf K},\bm{0}|\hat{G}(z){\cal H}|\beta,{\bf K},\bm{\delta}\rangle$ (13)
The second term can be expanded in terms of other propagators, leading to a linear system of coupled equations involving all these propagators. Although in principle the size of this system is infinite because $\bm{\delta}$ can be arbitrarily large, in practice the magnitude of the propagators decreases very fast [33] with increasing $\bm{\delta}$ for in-gap energies comparable to those of small excitons. This allows us to truncate this system by setting to zero all propagators with $|\bm{\delta}|$ larger than a cutoff, which is chosen such that the exciton energy converges within the desired accuracy. For very small excitons, good accuracy can be obtained even for a cutoff set to $2a$ (see example below) making such calculations extremely efficient. Full details and equations are given in Appendix B.
IV Results
IV.1 The range of validity of the continuum approximation
One way to validate the method described above is to verify that in the limit $U\rightarrow 0$, its results agree with those predicted by the continuum approximation. This will demonstrate that our method is also suitable to calculate the spectra of large excitons, although in this case the above-mentioned cutoff is large (it has to exceed the exciton radius) and the computation becomes much more expensive than for small excitons.
Figure 1: 1D exciton with K=0{K}=0 binding energy (left panel) and its radius in units $a=1$ (right panel) from the continuum approximation (blue line) and a 1D lattice model (red circles) with a one-orbital conduction band with $t_{c}=1$, and a $s,p_{x}$ two-orbital valence band with $t_{ss}=1.1,t_{pp}=-1,t_{sp}=0.5$. The corresponding valence band structure is shown in Fig. 2. Here, we assume $U_{s}=U_{p}=U$. The vertical dashed line marks the value of $U$ above which the continuum approximation starts to fail. See text for more details.
Figure 2: The two valence bands $E_{v}({k,\pm})$ (red lines) corresponding to the parameters used in Fig. 1. The top of the valence band is at $k=0$, as is the bottom of the conduction band (not shown), indicating a direct gap $K=0$. The dashed blue line shows the continuum approximation, which matches the exact dispersion to within $1\%$ for $|k|\lesssim\pi/6$ (dashed vertical lines).
Figure 1 compares the exciton binding energy for momentum K=0{K}=0, and the corresponding exciton radius $\xi$ as predicted by a 1D continuum model (blue lines; the relevant equations are listed in Appendix D) and a 1D lattice model with a $s,p_{x}$ two-orbital valence band (solid red circles). The magnitudes of the hopping integrals are $t_{c}=1$, $t_{ss}=1.1,t_{pp}=-1$ and $t_{sp}=0.5$ (the signs due to the appropriate orbital overlaps are included). For these values, the semiconductor has a direct gap, which is why we consider ${\bf K}=0$. This is demonstrated in Fig. 2, which shows the dispersions of the two valence bands $E_{v}({k},\pm)$ in the Brillouin zone (red lines).
For this specific model, the continuum approximation is accurate for $U\leq U_{c}\approx 2.2t_{c}$, after which it becomes increasingly worse, see Fig. 1. This corresponds to an exciton radius $\xi\gtrsim 2a$, which at first sight is a surprisingly low bound for a 'large' exciton. However, we can understand this as follows. For a given $\xi$, top-valence and conduction band states with momenta up to $k\sim 1/\xi$ contribute the most to the $K=0$ exciton $|K=0,{\rm exc}\rangle=\sum_{k}\phi_{k}c^{\dagger}_{k}v_{k,+}|{\rm GS}\rangle$. Here, $\xi\gtrsim 2a$ implies $|k|\lesssim{1\over 2a}\approx{\pi\over 6a}$. This value is shown by the vertical dashed lines in Fig. 2, and indeed provides a good bound for the top-valence band states described well by the continuum approximation (blue parabola).
We have verified for multiple other parameter choices (not shown) that we always get a good estimate for the smallest exciton radius $\xi\sim 1/k_{M}$ for which a continuum approximation is still accurate, when we choose $k_{M}$ as the largest momentum for which a quadratic approximation is still reasonable near the top/bottom of the valence/conduction bands of that specific model.
This simple procedure provides a practical recipe for estimating whether a continuum approximation is reasonable when the dispersions of the conduction and valence are known. As we point out in the next section, it may still be very misleading in a multi-orbital system.
IV.2 Qualitative failure of the continuum approximation in multi-orbital systems
In the example discussed above, we set $U_{s}=U_{p}$ for simplicity. However, in reality these quantities are highly dependent on the symmetry of the orbitals involved (their charge density) and their overlap with the conduction band orbital's charge density. In fact, there is no scenario that we can think of where $U_{s}=U_{p}$ is likely to hold.
Different Coulomb attractions for holes located in different orbitals add a second important ingredient in deciding the exciton spectra, besides the dispersion of the top valence band in the multi-orbital system, discussed above. This is because a stronger Coulomb attraction for one of the orbitals may lead to a larger binding energy for excitons involving valence band states which has a maximal character of that type, even if these states are not at the top of the valence band.
Figure 3: Gap energy $E_{\rm gap}(K)$ (black solid line) for the 1D lattice model with parameters as in Fig. 1. Red circles show the exciton dispersion $E_{\rm exc}(K)$ for $U_{p}=U_{s}=0.5$, while blue crosses show $E_{\rm exc}(K)$ for $U_{p}=3U_{s}=1.5$. The latter case stabilizes an exciton with momentum $K=\pi$ even though this is a direct gap ($K=0$) semiconductor, because the top-valence band states near $k=\pi$ have dominant $p$-character.
Such an example is exemplified in Fig. 3. The solid black line shows $E_{\rm gap}(K)$, i.e. the lower edge of the electron-hole continuum, for the same parameters as in Fig. 1. Clearly, this is a direct-gap semiconductor with the minimum gap at $K=0$. Accordingly, the continuum approximation always predicts that the maximum exciton binding energy must appear at $K=0$. The exciton spectrum of the 1D lattice agrees with this prediction if $U_{s}=U_{p}=0.5$, shown by the red circles. Indeed, here the strength of the attraction is the same independent of the character of the top-valence band, and therefore the most stable exciton appears where the electron-hole excitation gap is minimum. However, if we choose $U_{p}=3U_{s}=1.5$ (blue crosses), then the exciton minimum moves to $K=\pm\pi$ because the top-valence band has a predominant $p$-character near $k=\pi$ (for this model). Note that the exciton dispersion near $K=0$ is little affected by the increase of $U_{p}$, consistent with the fact that near $k=0$ the top-valence band has primarily $s$-character in our model.
To summarize, if the continuum approximation simplifies the multi-orbital valence bands' dispersion to a single parabola centered at the top of the top-valence band, it can never reproduce the phenomenology exemplified in this subsection, because information regarding the orbital character of the actual valence band is lost. Of course, one could proceed to a 'multi-valley' type of continuum approximation that would do better, however, that removes the top advantage of the continuum approximation, namely its ability to produce analytical results. If numerical computations are needed, studying a realistic lattice model becomes more efficient, beside being accurate when describing small excitons.
Figure 4: Heat maps showing the spectral weight $A_{\alpha\alpha}({\bf K},\omega)=-{1\over\pi}\text{Im}G_{\alpha\alpha}({\bf K},\bm{0},z)$ for $\bf K$} along high-symmetry lines in the Brillouin zone, when $\alpha=d_{x^{2}-y^{2}}$ (top panel), $\alpha=d_{xy}$ (middle panel) and $\alpha=d_{3z^{2}-r^{2}}$ (bottom panel). Here $U=12$ and all other parameters are as listed in Appendix A. The superimposed thick black line marks the lower edge of the electron-hole continuum, $E_{\rm gap}({\bf K})$. See text for more details.
IV.3 2D exciton transition
In this subsection, we illustrate another possible way for the continuum approximation to fail qualitatively by predicting the wrong momentum of the lowest energy exciton, in a multi-orbital model. Before proceeding, we mention that we verified that in the limit of weak electron-hole attraction, our method produces exciton results for this 2D model that are in agreement with those of the continuum approximation, see Appendices E-F.
As an example, we use the 2D triangular lattice with the three $d$ orbital bases introduced in Section II. Symmetry guarantees that their respective overlaps with an $s$-type conduction band orbital are the same, hence we set $U_{\alpha}=U$ for all three valence orbitals. (A short discussion of the effect of unequal $U_{\alpha}$ can be found in Appendix C). The other parameters are listed in Appendix A.
Figure 4 shows the spectral weight $A_{\alpha\alpha}({\bf K},\omega)=-{1\over\pi}\text{Im}G_{\alpha\alpha}({\bf K},\bm{0},z)$ for $\bf K$} along high-symmetry lines in the Brillouin zone, for $\alpha=d_{x^{2}-y^{2}}$ (top panel), $\alpha=d_{xy}$ (middle panel) and $\alpha=d_{3z^{2}-r^{2}}$ (bottom panel). The thick black line marks the lower edge of the electron-hole continuum, $E_{\rm gap}({\bf K})$. Above this edge, the spectral function contains continuum states, but on the color scale used, it is too weak to be clearly visible. The interesting features are the three exciton bands visible well inside the gap (due to the considerable $U$ value used). In the $d_{3z^{2}-r^{2}}$ projection (bottom panel) most of the intensity resides in the highest-energy branch and is therefore less relevant to the ground-state discussion. Along the $\Gamma-M-K$ cuts, the lowest branch appears with comparable intensity in the top two panels which show that it has comparable amounts of $d_{x^{2}-y^{2}}$ and $d_{xy}$ character for these momenta. By contrast, along the $K-\Gamma$ direction, top panel shows two exciton branches while the middle panel shows a third, different branch. Clearly, along this cut there is no mixing between these two orbitals, and the excitons have either $d_{x^{2}-y^{2}}$ or $d_{xy}$ character.
Figure 5: (top) Dispersion of the lowest-energy exciton $E_{\rm exc}({\bf K})$ for different values of $U$. The blue dashed line is the exciton energy calculated with a cut-off at $\delta=2$ (see Appendix G for details), and the dot marks its predicted lowest-energy exciton; (bottom) $E_{\rm exc}({\bf K})$ when ${\bf K}=M$ (blue triangles) and when $\bf K$} is the indirect gap, i.e. local minimum for $E_{\rm gap}({\bf K})$ along the $K-\Gamma$ line (red circles).
The continuum approximation would predict that the lowest energy exciton appears along the $K-\Gamma$ line, at the momentum $\bf K$} for which $E_{\rm gap}({\bf K})$ has a minimum; this is the momentum of the indirect gap of this semiconductor model. The top panel indicates that the lowest energy exciton branch has indeed a minimum at that momentum, however a more careful inspection shows that the global minimum is actually at the $M$ point. We note that the $M$ point is not even a local minimum for $E_{\rm gap}({\bf K})$!
Figure 6: Contour plots of the lowest-energy exciton band $E_{\rm exc}({\bf K})$ inside the Brillouin zone, for $U=2.5$ (top), $U=8$ (middle) and $U=12$ (bottom).
To understand what happens, we plot in the top panel of Fig. 5 the energy of the lowest-energy exciton branch along the same cuts, for increasing values of $U$. As expected, the exciton moves to lower energies as the attraction $U$ increases. Its dispersion follows closely the gap energy $E_{\rm gap}({\bf K})$ for small values of $U$, however, this changes as $U$ increases.
Figure 7: Probability $|\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle|^{2}$ plotted as a function of the distance $\bm{\delta}$ between electron and hole, for the lowest-energy exciton and for $\bf K$} equal to the indirect gap momentum, for $U=12$. The probabilities are normalized $\sum_{\beta,\bm{\delta}}|\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle|^{2}=1$. The projections are on orbital $\beta=d_{x^{2}-y^{2}}$ (top), $\beta=d_{xy}$ (middle) and $\beta=d_{3z^{2}-r^{2}}$ (bottom). The triangular lattice sites are shown by the white circles.
The bottom panel tracks the exciton energy at the momentum $M$ and at the indirect gap momentum (at minimum for $E_{\rm gap}({\bf K})$ along the $K-\Gamma$ line). The two curves cross at $U\approx 10$, marking a sharp transition in the momentum and character of the lowest-energy exciton. This is another example of a sharp transition that is impossible within the continuum approximation.
This change in the shape of the dispersion of the lowest-energy exciton band is illustrated graphically in Fig. 6, where the energy contour of $E_{\rm exc}({\bf K})$ is plotted in the Brillouin zone. For the smaller $U=2.5$ value (top panel) its minimum is clearly along the $\Gamma-K$ cut. However, as $U$ increases, the energy at the $M$ points drops faster such that for $U=12$ (bottom panel) the minimum is at the $M$ point.
Figure 8: Same as in Fig. 7 but for ${\bf K}=M$.
This transition can be attributed to the symmetry already mentioned, which prevents mixing of $d_{x^{2}-y^{2}}$ and $d_{xy}$ orbitals along the $\Gamma-K$ line. This is illustrated by the real-space probability plots to find the hole at a distance $\bm{\delta}$ from the electron in each of the three possible orbitals, shown in Fig. 7 for the indirect band-gap momentum and in Fig. 8 at the $M$ point.
First, both of these excitons are Frenkel-like, with most of the probability concentrated at $\bm{\delta}=0$. The exceptions are the $d_{x^{2}-y^{2}}$ and the $d_{3z^{2}-r^{2}}$ orbitals (top and bottom panel in Fig. 7, respectively) where symmetry prevents weight at $\bm{\delta}=0$ and instead there is a very small amount of weight at $|{\bm{\delta}}|=a$. This is consistent with the vanishing spectral weight for this exciton in the top plot of Fig. 4. By contrast, the $M$ point exciton has considerable on-site probability in all orbitals, see Fig. 8).
[…]
Finally, while the discussion was focused on the lowest-energy exciton and the transition with $U$ in its character and momentum, our results also show that as $U$ grows, additional exciton branches detach from the continuum, such that there are 3 branches for $U=12$. Their energies, orbital character and real-space wavefunctions (or probabilities) can be calculated similarly.
V Conclusions
In this work, we adapt a well-established approach for studying interacting few-fermion systems to the study of exciton spectra in multi-orbital lattice models. This method is particularly efficient for small excitons, such as those known to appear in a variety of semiconductors.
The results of the lattice models agree with the predictions of the continuum approximation in the limit of weak electron-hole attraction. This comparison allowed us to propose a simple criterion to estimate when the continuum approximation is likely to be accurate.
However, we also demonstrated that multi-orbital models exhibit exciton physics that is impossible to anticipate with a simple continuum approximation. Specifically, we presented two simple examples in which the momentum of the lowest-energy exciton does not match the momentum of the gap in certain regions of the parameter space. These examples only scratch the surface in terms of possible new behavior of small excitons in complex, multi-orbital bands, but help to illustrate the need to use realistic lattice models when modeling small excitons in specific semiconductors.
VI Acknowledgments
We acknowledge support from the Natural Sciences and Engineering Research Council of Canada (NSERC). We gratefully acknowledge the use of computing resources from the Stewart Blusson Quantum Matter Institute computing cluster LISA. Lastly, we also acknowledge that the research is done on the UBC Point Grey (Vancouver) campus, which sits on the traditional, ancestral, unceded territory of the $x^{w}m$$\theta k^{w}$$\overset{,}{y}$@$m$ (Musqueam) First Nation.
Appendix A Hoppings in the 2D, three-orbital valence band
We work in the orbital basis $(d_{x^{2}-y^{2}},\,d_{xy},\,d_{3z^{2}-r^{2}})$. We choose these $d$ orbitals as an example and for their non-trivial symmetry with respect to the underlying lattice; any other basis can be treated similarly. Following Harrison's rules and the Slater-Koster matrix elements [30, 34], the nearest-neighbor hopping integrals for a bond parallel to $\hat{x}$ are parameterized as follows: if the $d_{3z^{2}-r^{2}}\leftrightarrow d_{3z^{2}-r^{2}}$ hopping is denoted $t_{v3}$, then the $d_{x^{2}-y^{2}}\leftrightarrow d_{x^{2}-y^{2}}$ hopping is $t_{v1}/t_{v3}=19/9$, the $d_{xy}\leftrightarrow d_{xy}$ hopping is $t_{v2}/t_{v3}=-16/9$, and the $d_{x^{2}-y^{2}}\leftrightarrow d_{3z^{2}-r^{2}}$ hopping $t_{v4}/t_{v3}=-5\sqrt{3}/9$; all other hopping integrals are zero for symmetry reasons.
For example, $t_{v1}=\tfrac{3}{4}V_{dd\sigma}+\tfrac{1}{4}V_{dd\delta}$, $t_{v3}=\tfrac{1}{4}V_{dd\sigma}+\tfrac{3}{4}V_{dd\delta}$ so ${t_{v1}}/{t_{v3}}=(3r+1)/(r+3)$ where $r={V_{dd\sigma}}/{V_{dd\delta}}$ and $V_{dd\sigma}$ and $V_{dd\delta}$ are the two-center Slater–Koster $dd\sigma$ and $dd\delta$ matrix elements between neighboring $d$ orbitals. Using Harrison’s estimate $r=V_{dd\sigma}/V_{dd\delta}\approx(-6)/(-1)$ gives $t_{v1}/t_{v3}=19/9$. The other ratios are obtained similarly.
In matrix form, the hoppings along the bond $^{\bm{\delta}}_{0}=\hat{x}$ are, thus:
$t_{\alpha\beta}({\bm{\delta}_{0}})\;=\;\begin{pmatrix}t_{v1}&0&t_{v4}\\ 0&t_{v2}&0\\ t_{v4}&0&t_{v3}\end{pmatrix}$. (14)
The hopping matrix for a bond $\bm{\delta}_{b}$} oriented at an angle $\theta$ from $\hat{x}$ are generated by rotations in the $(x,y)$ plane:
$t_{\alpha\beta}({\bm{\delta}_{b}})\;=\;\underline{R}^{\!\top}(\theta)\,t_{\alpha\beta}({\bm{\delta}_{0}})\,\underline{R}(\theta).$ (15)
where
$\underline{R}(\theta)\;=\;\begin{pmatrix}\cos 2\theta&\;\sin 2\theta&0\\ -\sin 2\theta&\;\cos 2\theta&0\\ 0&0&1\end{pmatrix}$ (16)
In a triangular lattice, there are three inequivalent directions $\theta\in\{0^{\circ},60^{\circ},120^{\circ}\}$ corresponding to the three bond vectors $\bm{\delta}_{0}=(1,0)$, $\bm{\delta}_{1}=(\tfrac{1}{2},\tfrac{\sqrt{3}}{2})$, $\bm{\delta}_{2}=(-\tfrac{1}{2},\tfrac{\sqrt{3}}{2})$; the remaining three nearest-neighbors are their opposites $\bm{\delta}_{b}$ ($b=0,1,2$).
For the specific calculations shown in the main paper, we set $t_{v3}=1$, the electron hopping in the conduction band is $t_{c}=1$, and $\Delta=16.5$.
Appendix B Recurrence relations and their convergence
B.1 1D chain formalism
The 1D chain Hamiltonian with a multi-orbital valence band is given in Eqs. (2)–(6). We work in the exciton basis $|\alpha,K,\delta\rangle$, where $K$ is the center-of-mass crystal momentum, $\delta\in\mathbb{Z}$ is the electron–hole separation (in lattice spacings), and $\alpha\in\{s,p\}$ indexes the valence orbital. For a fixed source orbital $\gamma\in\{s,p\}$, we collect the two orbital components of the resolvent into the vector Green's function
$\bm{G}_{\gamma,\delta}(K,z)\equiv\begin{bmatrix}G_{\gamma,s}(K,\delta,z)\\[2.0pt] G_{\gamma,p}(K,\delta,z)\end{bmatrix}$.
We start by considering the action of the Hamiltonian $H$ on a ket. For $\delta\neq 0$, the on-site part of $H$ leaves $\delta$ unchanged, while hopping changes $\delta\to\delta\pm 1$.
For the $s$-component, Eq. (13) then becomes:
$(z-\Delta)\,G_{\gamma,s}(K,\delta,z)=$
$\quad\big(t_{ss}e^{-iK/2}-t_{c}e^{+iK/2}\big)\,G_{\gamma,s}(K,\delta{+}1,z)$
$\quad+t_{sp}e^{-iK/2}\,G_{\gamma,p}(K,\delta{+}1,z)$
$\quad+\big(t_{ss}e^{+iK/2}-t_{c}e^{-iK/2}\big)\,G_{\gamma,s}(K,\delta{-}1,z)$
$\quad+t_{sp}e^{+iK/2}\,G_{\gamma,p}(K,\delta{-}1,z),$ (17)
The equation for $G_{\gamma,p}(K,\delta,z)$ looks similar but $t_{ss}\to t_{pp}$.
We then group them in the recurrence relation:
$\underline{\alpha}\,\bm{G}_{\gamma,\delta}(K,z)\;=\;\underline{\beta}\,\bm{G}_{\delta+1}(K,z)\;+\;\underline{\gamma}\,\bm{G}_{\delta-1}(K,z),$ (18)
where for $\delta\neq 0$ the $2\times 2$ matrices are:
$\underline{\alpha}$ $=\begin{bmatrix}z-\Delta&0\\[2.0pt] 0&z-\Delta\end{bmatrix}$, (19)
$\underline{\beta}$ $=\begin{bmatrix}t_{ss}e^{-iK/2}-t_{c}e^{+iK/2}&t_{sp}e^{-iK/2}\\[2.0pt] t_{sp}e^{-iK/2}&t_{pp}e^{-iK/2}-t_{c}e^{+iK/2}\end{bmatrix}$, (20)
$\underline{\gamma}$ $=\begin{bmatrix}t_{ss}e^{+iK/2}-t_{c}e^{-iK/2}&t_{sp}e^{+iK/2}\\[2.0pt] t_{sp}e^{+iK/2}&t_{pp}e^{+iK/2}-t_{c}e^{-iK/2}\end{bmatrix}$. (21)
These are solved in terms of the ansatz $\bm{G}_{\gamma,\delta}(K,z)=\underline{S}_{\pm}\bm{G}_{\gamma,\delta\mp 1}(K,z)$, where the upper/lower signs correspond to $\delta>0$ and $\delta<0$, respectively, and $\underline{S}_{+}=[\underline{\alpha}-\underline{\beta}\underline{S}_{+}]^{-1}\underline{\gamma}$ and $\underline{S}_{-}=[\underline{\alpha}-\underline{\gamma}\underline{S}_{-}]^{-1}\underline{\beta}$ are continued fractions of matrices, calculated with $M$ iterations starting from ${\underline{S}}_{\pm}=0$. The cutoff $M$ is chosen sufficiently large to achieve convergence to the desired accuracy (see below). Physically, $M$ defines a variational solution whereby the electron and hole are not allowed to be $M$ or more sites apart.
The explicit form of Eq. (13) for $\delta=0$ reads:
$(\underline{\alpha}+\underline{\mathcal{U}})\,\bm{G}_{\gamma,0}\;=\;\mathbf{e}_{\gamma}\;+\;\underline{\beta}\,\bm{G}_{\gamma,1}\;+\;\underline{\gamma}\,\bm{G}_{\gamma,-1},$ (22)
where $\underline{\mathcal{U}}=\mathrm{diag}(U_{s},U_{p})$, $\mathbf{e}_{\gamma}=\begin{bmatrix}\delta_{\gamma s}\\[2.0pt] \delta_{\gamma p}\end{bmatrix}$. It is solved by using $\bm{G}_{\gamma,\pm}=\underline{S}_{\pm}\bm{G}_{\gamma,0}$ to find:
$\bm{G}_{\gamma,0}(K,0,z)\;=\;\big[\,\underline{\alpha}+\underline{\mathcal{U}}-\underline{\beta}\,\underline{S}_{+}-\underline{\gamma}\,\underline{S}_{-}\,\big]^{-1}\,\mathbf{e}_{\gamma}.$ (23)
From this we can then get any $\bm{G}_{\gamma,\delta}={\underline{S}}_{\pm}^{|\delta|}\bm{G}_{\gamma,0}$, depending on whether $\delta$ is positive or negative.
B.2 2D lattice formalism
The method described above can be generalized straightforwardly to lattices in higher dimensions [35]. Here we briefly describe the approach for a triangular lattice. The relative distance between electron and hole is $\bm{\delta}\in\mathbb{Z}^{2}$ and we define $N=|\bm{\delta}|$ (in units of $a$) as the minimum number of hops to get from an onsite configuration ($\delta=0$) to the relative separation $\bm{\delta}$. For a triangular lattice there are $6N$ distinct $\bm{\delta}$} separations for a given $N$, lying on a circular shell. We index these counter-clockwise starting from the one lying along the +x+x-axis, and group together into $\bm{G}_{\gamma,N}$ the $d_{N}=18N$ Green's functions with the same $N$. The extra factor of 3 is because we consider 3 possible valence orbitals per site.
In models with nearest-neighbor hopping, the equation of motion Eq. (13) for any Green's function with a given $N$ only links it to other Green's functions with the same $N$ or with $N\pm 1$. As a result, they can be again combined into a matrix recurrence relation, where for any $N\geq 1$:
$\underline{\alpha}_{N}(\mathbf{K},z)\,\bm{G}_{\gamma,N}=\underline{\beta}_{N}(\mathbf{K})\,\bm{G}_{\gamma,N+1}+\underline{\gamma}_{N}(\mathbf{K})\,\bm{G}_{\gamma,N-1},$ (24)
The matrices $\underline{\alpha}_{N}\in\mathbb{C}^{3d_{N}\times 3d_{N}}$, $\underline{\beta}_{N}\in\mathbb{C}^{3d_{N}\times 3d_{N+1}}$, $\underline{\gamma}_{N}\in\mathbb{C}^{3d_{N}\times 3d_{N-1}}$ can be read off from Eq. (13). Eq. (24) is solved using the ansatz $\bm{G}_{\gamma,N}={\underline{S}}_{N}(\mathbf{K},z)\bm{G}_{\gamma,N-1}$, where ${\underline{S}}_{N}(\mathbf{K},z)=[\underline{\alpha}_{N}(\mathbf{K},z)-\underline{\beta}_{N}(\mathbf{K}){\underline{S}}_{N+1}(\mathbf{K},z)]^{-1}\underline{\gamma}_{N}(\mathbf{K})$ is calculated iteratively starting from $\underline{S}_{M}(\mathbf{K},z)=0$ for a sufficiently large $M$ such that the desired accuracy is reached (see below). In particular, this allows us to calculate $\underline{S}_{1}(\mathbf{K},z)$.
The equation for $N=0$ reads:
$(\underline{\alpha}_{0}+\underline{\mathcal{U}})\bm{G}_{\gamma,0}$ $=\mathbf{e}_{\gamma}+\underline{\beta}_{0}(\mathbf{K})\,\bm{G}_{\gamma,1}$ (25)
where
$\underline{\mathcal{U}}$ $=\mathrm{diag}(U_{x^{2}-y^{2}},U_{xy},U_{3z^{2}-r^{2}})$ (26)
and the $\mathbf{e}_{\gamma}$ vector has an entry 1 for orbital $\gamma$, and 0 for the other two orbitals.
Figure 9: Relative error $\epsilon_{M}=|E_{\rm exc}(M)/E_{\rm final}-1|$ as a function $M$, where $E_{\text{exc}}(M)$ is the lowest exciton energy calculated using truncation at $M$, while $E_{\text{final}}$ is the converged value for the lowest exciton energy, obtained by truncating at a sufficiently large cutoff such that increasing it by 1 results in a relative change below $10^{-6}$. (a) 1D results: blue circles correspond to a large $U_{p}=U_{s}=2.2$, while red triangles correspond to a weaker $U_{p}=U_{s}=0.8$. In both cases $t_{c}=1,t_{ss}=1.1,t_{sp}=0.5,t_{pp}=-1.0$. The former case has a smaller exciton (the dashed vertical lines show the corresponding exciton radius $\xi$) and thus converges faster with increasing $M$; (b) corresponding 2D results for the three band model with $U_{x^{2}-y^{2}}=U_{xy}=U_{3z^{2}-r^{2}}=12$ and the other parameters listed in Appendix A. Here $\xi\approx 0.856$ (not shown) is very small and the convergence with M is very rapid.
The solution of Eq. (25) is, then:
$\bm{G}_{\gamma,0}(\mathbf{K},z)=\Big(\underline{\alpha}_{0}+\underline{\mathcal{U}}-\underline{\beta}_{0}\,{\underline{S}}_{1}(\mathbf{K},z)\Big)^{-1}\mathbf{e}_{\gamma}.$ (27)
Truncating at shell radius $N=1$ (i.e. setting $M=2$ when calculating the continued fractions $\underline{S}_{N}$), gives the one-shell approximation used in Appendix G. The cut-off used for Fig. 4-8 in the main text is $M=10$.
The convergence in terms of the cutoff $M$ is illustrated in Fig. 9 for sample 1D (top) and 2D (bottom) cases. The stronger the attraction and therefore the smaller the exciton, the faster the convergence with increasing $M$. For really small excitons, even a cutoff $M=2$ can already give reasonable accuracy, as further discussed in Appendix G. This illustrates the main advantage of using this method for strongly bound excitons.
Appendix C Longer-range Coulomb attraction
Inclusion of longer-range Coulomb interactions is trivial in this method: such terms are added to the corresponding $\underline{\alpha}$ matrix, similar to what was done above for $N=0$. As a result, the computational cost is the same as for an on-site attraction, if the resulting excitons have comparable radii.
Figure 10: Heat map similar to Fig. 4 showing the spectral weight $A_{\alpha\alpha}({\bf K},\omega)=-{1\over\pi}\text{Im}G_{\alpha\alpha}({\bf K},\bm{0},z)$ for $\bf K$} along high-symmetry lines in the Brillouin zone, summing over all the source orbitals. Parameters are: on-site attraction $U=11$, nearest-neighbor attraction $U_{1}=2$, all other parameters are as defined in Appendix A.
As an example, in Fig. 10 we show the exciton spectrum when a nearest neighbor attraction $U_{1}=2$ is included beside the on-site $U=11$; all other parameters are as listed in Appendix A. Importantly, we find that adding a nearest-neighbor attraction preserves the sharp ground-state momentum transition discussed in the main text, and in fact it provides additional tunability to achieve it. Arbitrarily long-range attraction can be added and studied similarly, however very small excitons only 'sample' short relative distances so whether there is attraction at distances larger than the cutoff $M$ is irrelevant.
Orbital-dependent on-site values $U_{\alpha}$ for the Coulomb attraction between the electron and the hole in orbital $\alpha$ can be considered similarly, by setting $\underline{\mathcal{U}}=\mathrm{diag}(U_{\alpha})$. In the main text we considered the case $U_{\alpha}=U$ to highlight the sharp transition in the ground state. Allowing orbital anisotropy $U_{\alpha}$ will change the results quantitatively, but the sharp transition persists except for highly fine-tuned parameters where the $U_{\alpha}$ vary very significantly with $\alpha$. An example plot of the exciton spectrum for $U_{x^{2}-y^{2}}=9.5,\,U_{xy}=10.5,\,U_{3z^{2}-r^{2}}=9.5$ is shown in Fig. 11.
Figure 11: Similar heat map as in Fig. 10, but now for on-site attraction $U_{x^{2}-y^{2}}=9.5,\,U_{xy}=10.5,\,U_{3z^{2}-r^{2}}=9.5$ and nearest-neighbor ${U}_{1}=0$.
Generalizations to orbital-dependent longer-range Coulomb attraction and/or inclusion of crystal field splittings for the valence orbitals can be treated similarly. We emphasize again that the computational cost is independent of such details, for comparable exciton radii.
Appendix D 1D Continuum approximation
1D continuum approximation replaces the lattice dispersion of the electron and hole with parabolic bands, as if the lattice constant $a\rightarrow 0$. The corresponding counterpart of the 1D Hamiltonian with on-site Coulomb attraction becomes:
$H=-\frac{\hbar^{2}}{2m_{e}}\frac{\partial^{2}}{\partial x_{e}^{2}}-\frac{\hbar^{2}}{2m_{h}}\frac{\partial^{2}}{\partial x_{h}^{2}}-Ua\delta(x_{e}-x_{h})$ (28)
where $x_{e},x_{h}$ are the (continuous variable) positions of the electron and hole along the chain, and $m_{e},m_{h}$ are their effective masses. Here $a$ is a characteristic length scale, which we identify with the lattice constant when comparing continuum and lattice predictions. For our 1D lattice model, the conduction band dispersion is $E_{c}(k)\!=\Delta\!-2t_{c}\cos(ka)$, corresponding to an effective electron mass $m_{e}={\hbar^{2}}/(2\,|t_{c}|\,a^{2})$ in the limit $k=0$. For the valence band, we diagonalize the multi-orbital model to find the dispersion of the top valence band $E_{v}(k,+)$, and extract the effective hole mass from its $k\rightarrow 0$ limit, $m_{h}=-\hbar^{2}/\frac{d^{2}E_{v}(k,+)}{dk^{2}}|_{k=0}.$
Hamiltonian (28) can be factorized into center-of-mass (CM) $X_{CM}=(m_{e}x_{e}+m_{h}x_{h})/M$ where $M={m_{e}+m_{h}}$, and relative $x=x_{e}-x_{h}$ components. CM motion with total momentum $K$ simply increases the total exciton energy by $\hbar^{2}K^{2}/2M$. The $K=0$ exciton binding energy $E_{\rm bind}>0$ is obtained from :
$\left(-\frac{\hbar^{2}}{2\mu}\frac{\partial^{2}}{\partial x^{2}}-Ua\delta(x)\right)\phi(x)=-E_{\rm bind}\phi(x),$ (29)
where $\mu=m_{e}m_{h}/M$ is the reduced mass. This gives:
$E_{\text{bind}}=\frac{\hbar^{2}}{2\mu\xi^{2}}=\frac{\mu(Ua)^{2}}{2\hbar^{2}}.$ (30)
where the exciton radius is $\xi=\kappa^{-1}=\frac{\hbar^{2}}{\mu Ua}$.
This can be generalized trivially to an indirect gap semiconductor: if the minimum gap corresponds to a momentum $K_{g}$, then the CM contribution to the exciton energy changes to $\hbar^{2}(K-K_{g})^{2}/2M$ while $E_{\text{bind}}$ is unaffected.
Appendix E 2D continuum approximation
Similar to the 1D case, the two–body continuum Hamiltonian separates into center–of–mass and relative parts, with the former contributing an additional $\hbar^{2}({\bm{K}}-{\bm{K}}_{g})^{2}/2M$ to the exciton energy. To calculate the exciton binding energy, we need to regularize the on-site attraction into a circular well of radius $R_{0}$ whose integrated strength is $g$:
$V(r)=-\,\frac{g}{\pi R_{0}^{2}}\Theta(R_{0}-r)$ (31)
where $\Theta$ is the Heaviside function. We solve the radial Schrödinger equation to find $\phi(r)\sim J_{0}(\kappa r)$ if $r<R_{0}$, and $\phi(r)\sim K_{0}(\kappa r)$ if $r>R_{0}$, where $J_{0},K_{0}$ are order-zero Bessel function of the first and second kind, respectively, and $\kappa=\sqrt{{2\mu E_{\rm bind}}/{\hbar^{2}}}$. Matching the radial wavefunction and its derivative at $r=R_{0}$, we find:
$E_{\rm bind}=\frac{\hbar^{2}}{2\mu R_{0}^{2}}\,\exp\!\Big[-\,\frac{4\pi\hbar^{2}}{\mu g}\Big].$ (32)
(Any additional pre-factor of order 1 from the matching can be absorbed into the microscopic length $R_{0}$, which acts as a fitting parameter of the binding energy.)
In the asymptotic limit $r\rightarrow\infty$, $K_{0}(\kappa r)\rightarrow{e^{-\kappa r}}/{\sqrt{r}}$ and we define the exciton radius as $\xi=1/\kappa$.
Appendix F Continuum vs. lattice results in 2D
Figure  12 shows a comparison between the lowest-energy exciton binding energy predicted by the 2D lattice model (red circles) and by its corresponding continuum model (blue solid line). The parameters for the lattice model are $t_{c}=1,t_{v1}=0,t_{v2}=0,t_{v3}=1,t_{v4}=0,\Delta=7$, but qualitatively similar results are obtained for any other choice of parameters. For the continuum model, we use the lattice results for weak attraction $U\in[0,2]$ to fit $R_{0}$ and $g$ from Eq. (32), and then generate the continuum prediction for all $U$.
Figure 12: Comparison between the lowest-exciton binding energy predicted by the 2D lattice model (red circles) and by the continuum model (blue line). Here, the parameters for the lattice model are $t_{c}=1,t_{v1}=0,t_{v2}=0,t_{v3}=1,t_{v4}=0,\Delta=7$, while those for the continuum model are obtained by fitting the lattice results with Eq. (32) for $U\in[0,2]$. The comparison is plotted (a) on a semi-log scale vs. $1/U$; and (b) on linear scale vs. $U$. As expected, the agreement is excellent for weak attraction $U$ but significant deviations appear for stronger attraction $U$. See text for more details.
The upper panel shows the comparison on a semi-log scale against $1/U$, and verifies that for weak attraction $U$, $\ln E_{\text{bind}}$ is indeed linear in $1/U$ as predicted by Eq. (32). However, the lower panel shows that on a linear scale, the two predictions start to disagree significantly for $U\gtrsim 6$ (for these parameters). This is consistent with the fact that for these lattice parameters, the bandwidth of the valence band is $W=9$. Indeed, disagreement with the continuum prediction is expected for $|U|\gtrsim W$ when all the valence states (not just the ones at the top of valence band, which are described well by the continuum model) are involved in the exciton wavefunction.
Appendix G Minimal variational approximation for the 2D strongly bound excitons
If the Coulomb attraction is so strong that the exciton is Frenkel-like, we expect a reasonable estimate to be obtained if we restrict the relative distance between electron and hole to be $|{\bm{\delta}}|\leq a$. One could then diagonalize the full Hamiltonian within this variational space, but the same results can be obtained efficiently from our method by setting all the propagators $\bm{G}_{\gamma,N}=0$ for $N\geq 2$, i.e. by choosing the cutoff at $M=2$.
In this case, the infinite recurrence relation for the vectors $\bm{G}_{\gamma,N}$ derived in Section B simplifies to:
$(\underline{\alpha}_{0}+\mathcal{U})\,\bm{G}_{\gamma,0}$ $=\mathbf{e}_{\gamma}+\underline{\beta}_{0}(\mathbf{K})\,\bm{G}_{\gamma,1},$
$\underline{\alpha}_{1}(\mathbf{K},z)\,\bm{G}_{\gamma,1}$ $=\underline{\gamma}_{1}(\mathbf{K})\,\bm{G}_{\gamma,0},$ (33)
so that $\bm{G}_{\gamma,0}=\left[\underline{\alpha}_{0}+\mathcal{U}-\underline{\beta}_{0}(\mathbf{K})[\underline{\alpha}_{1}(\mathbf{K},z)]^{-1}\underline{\gamma}_{1}(\mathbf{K})\right]^{-1}\mathbf{e}_{\gamma}$. The exciton energies are the poles of $\bm{G}_{\gamma,0}$ located inside the gap: $z<E_{\rm gap}({\bm{K}})$, and are found from:
$\det\!\Big[(z-\Delta+U)I_{3}-\;\underline{\beta}_{0}(\mathbf{K})\;\underline{\alpha}_{1}(\mathbf{K},z)^{-1}\;\underline{\beta}_{0}(\mathbf{K})^{\dagger}\Big]=0.$ (34)
Here, we assumed the same attraction $U_{\alpha}=U$ for all orbitals, and used $\underline{\gamma}_{1}(\mathbf{K})=\underline{\beta}_{0}(\mathbf{K})^{\dagger}$ which follows because the Hamiltonian is Hermitian.
As discussed in the main text (see blue dashed line Fig. 5), this very simple approximation indeed provides results with decent quantitative agreement at sufficiently large $U$. Of course, if the exciton is spread over more sites, the agreement is improved by increasing $M$ accordingly. Clearly, the results must converge eventually if $M$ is sufficiently large – this is the essence of the fully converged solution described in Section B.